The ratio of yes votes to no votes in an election was 8 to 5. How many total votes were cast if there were 4624 yes votes cast?

Answer by actuary(112)

Since the ratio of "yes" to "no" votes is 8 to 5, the number of "no" votes is found by solving the proportion 8/5 = 4624/n for n.
n = 4624*5/8= 2890.
The total votes cast is the sum of the
"yes" votes and the "no" votes = 4624 + 2890 = 7514
I hope that this helps.